OpenStudy (anonymous):

tan^2ø(1-sin^2ø)=sin^2ø

OpenStudy (anonymous):

\[\tan ^{2}\emptyset \left( 1-\sin^{2}\emptyset\right) = \sin ^{2}\emptyset \]

OpenStudy (anonymous):

\[(1-\sin^2(\theta)) = \cos^2(\theta)\]\[\tan^2(\theta)*(\cos^2(\theta)) = \sin^2(\theta)\]\[(\sin^2(\theta)/(\cos^2(\theta)))*\cos^2(\theta) = \sin^2 (\theta)\]\[\sin^2(\theta) = \sin^2(\theta)\]

OpenStudy (anonymous):

thank you :D